QUESTION 11 Consider the following information n1 100 , χ!= 60, n2= 100, and X2=40. The statistics analyst wants to know if there is evidence of a significant differences between tre two population proportions?Refer to the information given on question 11 the p-value for a two-tailed test is 1. 0.0023 2. 0.9977 3. 0.0046 4. 0.9974 5. 0.0026
Answer #1

Please read the answer below. Don't hesitate to give a "thumbs up" in case you're satisfied with the answer.

p = x1+x2/n1+n2 = 60+40/100+100 = .50

p1 = x1/n1 = 60/100 = .60

p2 = x2/n2 = 40/100 = .40

z =((0.6-0.4)-(0))/SQRT(0.5*0.5*(1/100+1/100)) = 2.828

The p-value is therefore, P(|z|>2.828) = 2*(1-.002339) = .004678

Answer is 3. 0.0046